OpenStudy (anonymous):

What is the value of the expression |20 + 5|?

OpenStudy (anonymous):

@dan815 @bibby

OpenStudy (anonymous):

25

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

@mathkid10 can you help me with this? explain how you find the answer please |4 + x| 12 What is the value of the expression when x = 6?

OpenStudy (shamil98):

is there no sign after the absolute value?

OpenStudy (anonymous):

No

OpenStudy (sleepyhead314):

do it like you would 12(4+6) except add the insides first

OpenStudy (anonymous):

22?

OpenStudy (anonymous):

That is not an option choice though

OpenStudy (sleepyhead314):

ehhh number(number) = multiply them

OpenStudy (sleepyhead314):

so 12 times (4+6) = ?

OpenStudy (anonymous):

its a -12 though so do i multiply it like a negative?

OpenStudy (shamil98):

|x+4| - 12 = |6+4| - 12 = do the rest

OpenStudy (anonymous):

i get -10 if i do it that way

OpenStudy (anonymous):

why does the - go to a +

OpenStudy (sleepyhead314):

|6+4| = |10| right? then what is 10 - 12 = ???

OpenStudy (anonymous):

-2

OpenStudy (sleepyhead314):

there you go

OpenStudy (anonymous):

thanks :)